The impulse response of a discrete-time LTI system is
h(n) = 2δ(n) − δ(n − 4).
Find and sketch the output of this system when the input is the step function
x(n) = u(n).
1
Expert's answer
2020-12-22T05:03:51-0500

"x(n) = u(n)."

"y(n) =x(n)\times [2\delta(n)-\delta(x-4)]"

"y(n) = 2u(n) - u(x-4)"